OpenStudy (anonymous):

Breann solved the system of equations below. What mistake did she make in her work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 20 x = 4 2(4) + y = 5 y = −3

OpenStudy (anonymous):

i don't see any mistake the answer is correct

OpenStudy (anonymous):

you can check that x = 4, y =-3 works in both equations \[ 2x+y=5\] \[2\times 4+(3)=5\] \[8-3=5\] true \[x-2y=10\] \[4-2(-3)=10\] \[4+6=10\] true

OpenStudy (anonymous):

thankyou , what about Systems of equations with the same slopes and different y-intercepts have no solution.

OpenStudy (anonymous):

another idiot math teacher question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

answers : always , never , or sometimes ?

OpenStudy (anonymous):

same slope an different y intercepts means lines are parallel

OpenStudy (anonymous):

Systems of equations with the same slopes and different y-intercepts have no solution. always jesus who makes up these questions????

OpenStudy (anonymous):

haha i know rihgttt